20 is what percent of 1480
Answer by stanbon(75887)

20 is what percent of 1480
----
x/100 = 20/1480
------
x = 100(20/1480)
---
x = 2000/1480
x = 1.35
----
20 is 1.35% of 1480
===============
Cheers,
Stan H.
===============